Which has a larger resistance, a 60 watt lightbulb or a 100 watt lightbulb ?
Answer #1

P=I*V
If V=230 V
for 100 W bulb,
100=230*I
I=100/230 = 0.43 A

Similarly for 60 W bulb
60=230*I
I=60/230 = 0.26 A

Low current High resistance.
Means 60 W bulb have more resistance